Problem 8.72

A uniform disk turns at 2.6 rev/s around a frictionless central axis. A nonrotating rod, of the same mass as the disk and length equal to the disk's diameter, is dropped onto the freely spinning disk . They then turn around the spindle with their centers superposed.

Part A What is the angular frequency in rev/s of the combination? Express your answer using two significant figures.
Answer #1

here angular momentum remain conserve Li = Lf

L = I*w

initially only disk is rotating

Ifor disk = 1/2*MR^2

Li = 2.6 * MR^2/2

Lf = (Irod + Idisk)wf

Irod = ML^2/12

L = 2R = MR^2/3

Lf = 5MR^2/6 *wf

Li = Lf

wf = 3*2.6/5

wf = 1.56 rev/s